Analyse the cross price elasticity of demand for wheat and rice when a change in the

price of wheat from Rs.70 to Rs. 90 results in a change in the quantity demand for

wheat from kg 3000 to kg 5000 in the market. Interpret the value of the coefficient.


1
Expert's answer
2021-04-22T07:50:30-0400

Cross-price elasticity=


"\frac{Percentage change in quantity of goodA}{percentage change in price of good B}"


Percentage change in quantity demanded= "\frac{5000-3000}{3000}×100=66.67%"%

Percentage change in price="\frac{90-70}{70}×100=28.57"%


Cross-price elasticity="\frac{66.67}{28.57}=2.33"


Cross price elasticity of wheat and rice is positive,thus they are substitute goods.